# Describe Where You Are: Improving Noise-Robustness for Speech Emotion Recognition with Text Description of the Environment

### v1

@@ 5.3 Evaluation of Different Types of Environmental Embedding
Our proposed method uses the embedding extracted from the text encoder to represent the testing environmental condition. To verify the benefits of using a text-based environmental embedding, we compare it with three different types of environmental embedding: one-hot encoding (One-hot), global vectors for word representation (GloVe) [47], and audio spectrogram transformer representation (AST) [48]. One-hot uses 20-dimension binary vectors, where 1 represents the target environment condition, and 0 represents the others. Each dimension corresponds to the environmental condition of the training set. This embedding fully represents a seen environment with a simple vector; however, it cannot represent unseen environments, which is inappropriate for real-world services. GloVe is a word-level vector representation extracted from the regression model that considers the co-occurrences of words. We import the pre-trained GloVe vector collections consisting of 2.2 million vocabularies. We select the word vector representation that corresponds to the target noisy environment. The resulting representation is a 300-dimension vector. This representation can deal with unseen environments using text description, but it is semantically limited compared to our proposed text encoders. AST uses a transformer architecture to map the spectrogram patches into an audio-level representation. The model is fine-tuned with sound event classification tasks by using AudioSet, which is the same noise sound corpus for our training set. This model can automatically capture the acoustic characteristics from the audio-only input. However, it cannot explicitly use the semantic information of the testing environment.
@@
We compare our proposed method with One-hot in the seen environment scenario (Table 2) and with the other baselines in the unseen environment scenario (Table 3). For the seen environment scenario, we used the same environmental condition as the train set to contaminate the clean test set but with different audio samples. We use ten different test sets and report the average CCC for both cases. Table 2 and 3 report the results for the seen and unseen environments, respectively. In the seen environment, our proposed method and the one-hot environment encoding model improve the original SER performance for all the conditions and attributes. Both models show similar performances in the seen environments. However, the one-hot encoding cannot cover unseen environments. This result demonstrates that the proposed text embedding can deal with both seen and unseen environments. Compared to the model that uses GloVe embeddings, our proposed method shows better SER performances in 0dB and -5dB conditions. It also shows a better performance for valence in the 10dB condition. The GloVe model only considers word co-occurrence to get a word embedding, while our proposed text encoder model is pre-trained to understand the semantic information of a sentence. This result implies the importance of pre-training the text encoder with language modeling to get a robust environment embedding for performance improvement. The AST strategy shows better performance for valence than our proposed model under the 5dB and 0dB conditions. In comparison, our proposed model performs better for all the emotional attributes under the -5dB conditions. AST does not use semantic information from the testing environment to get environmental embedding; instead, it extracts the environmental information from the given audio. Considering that the -5dB SNR level is not presented while training the model, the result demonstrates that the AST works well for the seen SNR level but not for the unseen SNR level. In contrast, our proposed method works well for the unseen SNR level since the text description is independent of the SNR level. This result demonstrates that our proposed method is robust against the unseen SNR level, which is practical for real-world scenarios.
@@ 5.4 Benefit of Fine-Tuning the Text Encoder
Table 4 reports the average CCC of ten different test sets for each model. When comparing the TG-EAT-LLM and TG-EAT-LLM-FT implementations, they do not show significantly different performance. However, the TG-EAT-CL-FT approach shows meaningful performance improvement over the TG-EAT-CL implementation. For the -5dB conditions, it even reaches the best performance among all the models. This observation illustrates the importance of compensating the embedding space gap between the pre-trained text encoder space and the acoustic embedding. Although jointly fine-tuning the text encoder and the SER model can cost more memory space and computation time for the adaptation, this strategy can fully utilize the potential of the text encoder pre-trained with the audio modality.
@@ 6 Conclusions
We proposed the TG-EAT method, which uses a text description of the testing environment for noise-robust SER. This approach inserts a text-based environment representation into an SER model, leading it to denoise the speech representation with respect to the given environmental information. Our experiment demonstrated that the LLM-based representation can improve SER performance under noisy conditions, especially when dealing with low SNR conditions. Our analysis indicates that the pre-trained text encoder can cluster acoustically and semantically similar environments into the same embedding, which is crucial for generalizing the models for unseen environments. Our result also shows that the CLAP-based text encoder can be highly improved by updating the text encoder. This result demonstrates the importance of minimizing the embedding space gap between the text encoder and the acoustic embedding.
We plan to expand this approach to cases where we cannot obtain information on the testing environment. We assume that the CL-based representation can address the scenario when the noise information is not provided by introducing its audio encoder. CLAP trains the audio encoder to have a similar representation to the ones from the text encoder, which could be useful for extracting environmental information from the audio. For this reason, we plan to investigate how we can improve the noise-robustness of the SER model with a CLAP encoder.

### v2

@@ 5.3 Evaluation of Different Types of Environmental Embedding
Our proposed method uses the embedding extracted from the text encoder to represent the testing environmental condition. To verify the benefits of using a text-based environmental embedding, we compare it with three different types of environmental embedding: one-hot encoding (One-hot), global vectors for word representation (GloVe) [Pennington_2014], and audio spectrogram transformer representation (AST) [Gong_2021]. One-hot uses 20-dimensional binary vectors, where 1 represents the target environment condition, and 0 represents the others. Each dimension corresponds to the environmental condition of the training set. This embedding fully represents a seen environment with a simple vector; however, it cannot represent unseen environments, which is inappropriate for real-world services. GloVe is a word-level vector representation extracted from the regression model that considers the co-occurrences of words. We import the pre-trained GloVe vector collections, which consist of a 2.2 million-word vocabulary. We select the word vector representation that corresponds to the target noisy environment. The resulting representation is a 300-dimensional vector. This representation can handle unseen environments through text description, but it is semantically limited compared to our proposed text encoders. AST uses a transformer architecture to map the spectrogram patches into an audio-level representation. The model is fine-tuned with sound event classification tasks using AudioSet, which serves as the noise sound corpus for our training set. We directly import the pre-trained checkpoint from HuggingFace and extract the patch-wise embedding sequence from the given input. We apply average pooling to the extracted sequence to yield a single environment embedding, which is then fused with the pre-trained SER model. We do not fine-tune the pre-trained checkpoint jointly with the SER model, following the same strategy we use to train the TG-EAT-LLM framework. This model can automatically capture the acoustic characteristics from the audio-only input. However, it cannot explicitly use the semantic information of the testing environment.
@@
We compare our proposed method with the one-hot vector in the seen environment scenario (Table 5) and with the other baselines in the unseen environment scenario (Table 6). For the seen environment scenario, we used the same environmental conditions as the train set to contaminate the clean test set, but with different audio samples. We use ten different test sets and report the average CCC for both cases. Tables 5 and 6 report the results for the seen and unseen environments, respectively. In the seen environment, our proposed method and the one-hot environment encoding model improve the original SER performance for all the conditions and attributes. Both models show similar performances in the seen environments. However, the one-hot encoding cannot cover unseen environments. This result demonstrates that the proposed text embedding can deal with both seen and unseen environments. Compared to the model that uses GloVe embeddings, our proposed method shows better SER performances in the 0dB and -5dB conditions. It also shows a better performance for valence in the 5dB condition. The GloVe model only considers word co-occurrence to get a word embedding, while our proposed text encoder model is pre-trained to understand the semantic information of a sentence. This result implies the importance of pre-training the text encoder with language modeling to get a robust environment embedding for performance improvement. The AST strategy significantly improves the performance for arousal and dominance. However, it fails to improve the performance for valence when the SNR level is high (e.g., 5dB and 0dB conditions). AST does not use semantic information from the testing environment to get environmental embedding; instead, it extracts the environmental information from the given audio. We hypothesize that AST confuses the environmental condition when the background noise amplitude is comparably lower than the speech sound. Unlike this approach, our proposed model relies on the text description, which is independent of the SNR level. Therefore, it performs better than AST for valence. In the 0 dB and -5 dB conditions, our method significantly improves the original models’ performances for all the attributes. Considering that those low SNR levels are not presented while training the model, the result demonstrates that our proposed method is robust against unseen SNR levels, which is practical for real-world scenarios.
((a))
((b))
Figure 4: Comparison of the text embedding projection obtained before fine-tuning (TG-EAT-CL) and after fine-tuning (TG-EAT-CL-FT). Similar to the plots in Figure 3, we use UMAP to project text embeddings into a 2D space.
(a) Arousal
(b) Dominance
(c) Valence
Figure 5: Relative improvement of fine-tuning the text encoder (TG-EAT-CL-FT) in the TG-EAT-CL framework under the -5dB condition. We illustrate 16 environments, including the eight highest and the eight lowest improvements for each attribute.
@@ 5.4 Benefit of Fine-Tuning the Text Encoder
Table 7 reports the average CCC of ten different test sets for each model. When comparing the TG-EAT-LLM and TG-EAT-LLM-FT implementations, they do not show significantly different performance. However, the TG-EAT-CL-FT approach shows meaningful performance improvement over the TG-EAT-CL implementation. For the -5dB conditions, it even reaches the best performance among all the models. When compared with TG-EAT-CL, fine-tuning the text encoder improves the recognition performance by 76.4% (arousal), 100.0% (dominance), and 27.7% (valence). To analyze how fine-tuning benefits the model, we visualize in Figure 4 the text embedding projection from the text encoder used for the TG-EAT-CL and TG-EAT-CL-FT models. We use the same experiment setting as we used for illustrating Figure 3. We can see that some of the embeddings that were not clustered well in TG-EAT-CL are corrected in TG-EAT-CL-FT. For example, “cafe” is distant from other indoor environments in TG-EAT-CL. However, when fine-tuning the text encoder, its embedding gets closer to those environments. We also observe that such cluster alignments could lead to performance improvement. Figure 5 illustrates the relative performance improvement of TG-EAT-CL-FT for each environmental condition. We can see that the TG-EAT-CL-FT model improves performance in the “cafe” environments. We can also see that the fine-tuning strategy can improve the performance for all the attributes, except for five environments in valence (“creek”, “playground”, “stadium”, “arcade”, and “fountain”). This observation illustrates the importance of compensating for the gap in the embedding space between the pre-trained text encoder space and the acoustic embedding. Although jointly fine-tuning the text encoder and the SER model can cost more memory space and computation time for the adaptation, this strategy can fully utilize the potential of the text encoder pre-trained with the audio modality.
@@
Table 8: Average CCC of the six sessions with the clean and noisy version of the MSP-IMPROV corpus. We compared the proposed method with the baselines by using the wavlm-base-plus model.
SNR Model Arousal Dominance Valence
Clean
Original 0.38 0.44 0.41
RT 0.39 0.44 0.40
TG-EAT-CL 0.38 0.45 0.44
TG-EAT-LLM 0.40 0.45 0.42
Random
Original 0.32 0.36 0.25
RT 0.40 0.42 0.30
TG-EAT-CL 0.36 0.42 0.34
TG-EAT-LLM 0.40 0.42 0.32
5.5 Cross-corpus Generalization
To evaluate the generalization ability of our proposed TG-EAT models under unseen, out-of-domain dataset, we conduct a cross-corpus analysis. We test our baselines and the proposed TG-EAT models on the MSP-IMPROV dataset [Busso_2017], where their data acquisition process is different from our training set, the MSP-Podcast Corpus. For this evaluation, we evaluate the performance for each of the six sessions in the MSP-IMPROV corpus. We create a noisy version of this test set by contaminating the clean audio with noise from the DEMAND database at a randomly selected SNR level between -5dB and 5dB.
Table 8 illustrates our experiment results. Under the clean condition, we can see that our proposed TG-EAT models do not degrade the performance of baselines. Indeed, they outperform the RT baseline for valence. As we discussed in Section 5.1, training with multiple conditions could help performance improvement for valence even under the clean condition. We can see that guiding the model with environmental conditions can keep this benefit under a cross-corpus scenario, while preserving the Original SER model’s generalization ability.
As expected, under noisy conditions, all models retrained with noisy speech (RT, TG-EAT-CL, and TG-EAT-LLM) significantly outperform the Original model, confirming that noise-aware training is crucial for robust performance on out-of-set corpora. Our models show a clear improvement in valence. As discussed in Section 5.1, arousal and dominance are more closely related to acoustic characteristics, while valence is correlated with linguistic content. By providing explicit environmental information, our TG-EAT framework may allow the SER model to better normalize for acoustic variability, thereby improving the extraction of linguistic content critical for valence prediction.
Interestingly, the TG-EAT-CL model yields the best valence performance under noisy, out-of-set conditions, despite not being the top performer in the in-set evaluation. This result indicates that its learned representation possesses a strong capability for generalization, particularly under challenging mismatched conditions. While our framework shows clear benefits, we note that the significant improvements observed for arousal and dominance in the in-set condition did not fully transfer to this cross-corpus task. This suggests that future work could explore methods to further enhance the generalization of acoustically-related emotion attributes in out-of-domain scenarios.
5.6 Impact of Mislabeled Audio Tags
In a real-world scenario, location tags could be mislabeled due to an inaccurate GPS signal or ambiguous locations. To evaluate the impact of mislabeling audio tags, we present an ablation study to report the performance of the proposed TG-EAT-LLM framework, manipulating the environmental description. To simulate mislabeling, we intentionally manipulated the environmental tags associated with the input noisy speech at varying levels of distortion. Specifically, the environmental tags were randomly replaced with ratios of 25%, 50%, and 75%. The original model trained without any mislabeled tags (0% manipulation) was used as the baseline for comparison.
Figure 6: CCC of the TG-EAT-LLM model trained with noisy speech paired with mislabeled environment conditions.
To ensure a robust evaluation, we tested the model on 10 different testing sets, each contaminated with noise at a randomly selected SNR between 5dB, 0dB, and -5 dB. The environmental conditions in the testing data were also chosen randomly to simulate diverse real-world scenarios. The average CCC across these 10 sets was computed for each manipulation ratio. Figure 6 illustrates the result of the ablation study. The performance of the TG-EAT-LLM model gradually decreases as the levels of mislabeled audio tags increase. The trend indicates that as the model is exposed to higher degrees of mislabeling, it struggles to cluster recordings accurately from similar environmental conditions. This performance degradation highlights the sensitivity of the TG-EAT-LLM framework to the quality of audio tags and the need for accurate labeling during training. We can see that environmental conditions play a significant role in our proposed framework. The model relies heavily on this information to achieve robust SER performance, highlighting the importance of minimizing labeling errors when describing environmental conditions.
Table 9: Silhouette score of the last hidden layer’s embedding space of the RT and TG-EAT-LLM models.
RT TG-EAT-LLM
Using the correct environmental tag 0.363 0.133
Using a semantically similar but incorrect tag 0.363 0.133
Using a semantically different tag 0.360 0.140
Using without an environmental tag 0.361 0.138
We investigate the degree of impact on our proposed model of using an audio tag that is mislabeled but either semantically similar or completely dissimilar to the correct environmental information. We measured the separability between the final layer embeddings of speech under clean versus noisy conditions. A robust model should normalize the environmental difference, which should lead to a low clustering quality when the embedding space is clustered by environmental differences. We compared our TG-EAT-LLM model against the RT baseline under four prompt conditions: (1) the correct environmental tag, (2) a semantically similar but incorrect tag, (3) a semantically different tag, and (4) no tag at all. We select five similar conditions (“kitchen”, “house”, “living room”, “school”, “office”), contaminating the test set with these noise types. SNR levels are randomly chosen from -5dB to 5dB. Condition (1) uses the same tag as the noise label in the input audio. Condition (2) randomly selects the tags from the four other similar conditions. Condition (3) randomly selects tags from five different tags that are semantically different from this group (“playground”, “subway station”, “town square”, “construction site”, “sports field”). Condition (4) does not use any environmental tags (i.e., the model only accepts the input audio).
Table 9 shows the result of our experiment. The TG-EAT-LLM model consistently achieves much lower silhouette scores than the RT baseline. When using a correct tag, the RT baseline yields a 0.363 score, while our TG-EAT-LLM’s score achieves a 0.133 score. This result illustrates that our framework effectively normalizes environmental differences from the speech representation. When compared to using the correct tag, using a semantically similar tag does not change the clustering score in TG-EAT-LLM. The score slightly increases when a semantically different tag or no tag is provided. The RT baseline’s score remains unchanged regardless of the text input. These observations imply that guiding the model with environmental information during training can introduce sensitivity to semantically incorrect environmental information. The model does not ignore the prompt but instead uses it as intended to disentangle environmental noise from emotional content.
5.7 Limitations
Our proposed TG-EAT framework heavily relies on the assumption that the recorded speech is paired with accurate GPS location data, which is crucial for acquiring accurate environmental tags. However, the recorded speech could be associated with inaccurate or missing GPS points in real-world scenarios, leading to irrelevant or unavailable environmental tags. As discussed in Section 5.6, having irrelevant tags could degrade our system’s performance, and missing tags would not provide the information for our model to work properly. Additionally, GIS mashups may fail to retrieve meaningful tags in areas with sparse or incomplete annotations, further limiting the system’s ability to leverage the information of the recording conditions. Furthermore, even with accurate data, a single static tag may be an oversimplification for complex acoustic scenes with overlapping speech or rapidly changing soundscapes, potentially degrading performance. Those limitations demonstrate the need for future work to address potential inaccuracies in GPS data and missing GPS modality. In addition to the availability of accurate GPS information, our architectural approach to fuse the text embedding to the SER model was limited to concatenating a text embedding to audio tokens. The exploration of more dynamic fusion strategies [Perez_2018, Alayrac_2022] remains a key area for future work to potentially build upon our findings and further enhance performance.
@@ 6 Conclusions
We proposed the TG-EAT method, which uses a text description of the testing environment for noise-robust SER. This approach inserts a text-based environment representation into an SER model, leading it to improve the prediction with respect to the given environmental information. Our experiment demonstrated that the LLM-based representation can improve SER performance under noisy conditions, especially when dealing with low SNR conditions. Our analysis indicates that the pre-trained text encoder can cluster acoustically and semantically similar environments into the same embedding, which is crucial for generalizing the models for unseen environments. Our result also shows that the CLAP-based text encoder can be highly improved by updating the text encoder. This result demonstrates the importance of minimizing the embedding space gap between the text encoder and the acoustic embedding.
We plan to expand this approach to cases where we cannot obtain information on the testing environment. While AST embeddings demonstrate competitive performance for arousal and dominance, they do not show improvements for valence compared to models that explicitly use text embeddings. The CL-based representation can address scenarios where noise information is not provided by introducing its audio encoder. CLAP trains the audio encoder to have a similar representation to the ones from the text encoder, which could be useful for extracting environmental information from the audio. For this reason, we plan to investigate how we can improve the noise-robustness of the SER model with a CLAP encoder. We also plan to investigate the alternative approach of leveraging the inferred caption from an automated audio captioning (AAC) model [Mei_2022], using it as an environmental descriptor